# Symmetry Breaking in Chemical Systems: Engineering Complexity through Self-Organization and Marangoni Flows

### v3

@@ Abstract
Far from equilibrium, chemical and biological systems can form complex patterns and waves through reaction-diffusion coupling. Fluid motion often tends to disrupt these self-organized concentration patterns. In this study, we investigate the influence of Marangoni-driven flows inside a thin layer of fluid ascending the outer surfaces of hydrophilic obstacles on the spatio-temporal dynamics of chemical waves in the modified Belousov-Zhabotinsky reaction. Our observations reveal that circular waves originate nearly simultaneously at the obstacles and propagate outward. In a covered setup, where evaporation is minimal, the wavefronts maintain their circular shape. However, in an uncovered setup with significant evaporative cooling, the interplay between surface tension-driven Marangoni flows and gravity destabilizes the wavefronts, creating distinctive flower-like patterns around the obstacles. Our experiments further show that the number of petals formed increases linearly with the obstacle’s diameter, though a minimum diameter is required for these instabilities to appear. These findings demonstrate the potential to ’engineer’ specific wave patterns, offering a method to control and direct reaction dynamics. This capability is especially important for developing microfluidic devices requiring precise control over chemical wave propagation.
@@
Sangram Gore, Binaya Paudyal, Mohamed Ali, Nader Masmoudi, Albert Bae, Oliver Steinbock, Azam Gholami*
@@
Lewis &\&& Clark College, Portland, Oregon, USA
@@ 1 Introduction
The interaction between chemical reactions and convective fluid motion, particularly through mechanisms such as Marangoni-driven flows, profoundly influences the formation and breakdown of complex spatio-temporal patterns [26, 27, 28, 29]. In pioneering studies on the BZ reaction [30, 31, 32, 33, 34], it was observed that in confined environments, chemical instabilities manifest themselves as propagating waves. In contrast, in systems with a free surface, the presence of Marangoni flows lead instead to the formation of cellular structures [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 12]. These flows are initiated by gradients in surface tension [46, 44], which might arise from either temperature variations or differences in solute concentrations across the fluid’s surface, causing the fluid to flow from regions of lower to higher surface tension. For instance, in the ferroin-catalyzed BZ reaction, Fe2+ complex has a higher surface activity than Fe3+ complex which is attributed to the differing hydrophilic properties of the catalyst’s oxidized (more hydrophilic) and reduced forms. This difference leads to an increase in surface tension of approximately 2 mN/m when the system transitions from the reduced to the oxidized state [46]. In an uncovered setup, where evaporative cooling is significant, the interplay of chemical waves, buoyancy-driven flows due to density gradients, and Marangoni flows due to surface tension gradients collectively influences the resulting chemo-hydrodynamic patterns [31, 47, 48, 41, 40, 49, 34, 50].
@@
Evaporation-induced Marangoni flows can be intensified by introducing hydrophilic vertical or inclined obstacles within the reactive medium. The thin film rising along the outer surfaces of these obstacles evaporates more quickly than the bulk fluid, creating thermal/solutal surface tension gradients. The delicate balance between upward-driving surface tension forces and gravity-induced downward motion leads to undulations at the fluid interface, potentially developing into finger-like patterns if the fluid becomes sufficiently unstable [51, 52, 53, 54, 55, 56, 57]. The obstacle’s geometry also affects the dynamics at the contact line region of thin films. A cylindrical obstacle introduces unique conditions, such as azimuthal curvature and periodic boundary conditions, absent in flat obstacles [57]. This periodicity ensures that an integer number of fingers can form along a cylinder, a condition that is not required for a flat obstacle, and the smallest number of fingers that can form along the contact line of a cylinder is one [57].
@@
Building on these groundworks, our study presents a novel approach to explore the intricate interaction between evaporation driven Marangoni flows and chemical waves in the modified BZ reaction in the presence of hydrophilic obstacles. We observe that in a covered experimental setup with minimal evaporative cooling, the obstacles act as focal points that consistently emit synchronous circular waves and maintain their shape over multiple cycles. In contrast, thermal/solutal driven Marangoni flows in an uncovered setup, where evaporative cooling is more pronounced, trigger a fingering instability that leads to the formation of distinct flower-like wave patterns. Significantly, the number of petals increases with the diameter of the obstacles, and a minimum threshold value for the diameter is required for this instability to occur. We performed numerical simulations of a reaction-diffusion model to explain the observed wave dynamics and gain deeper insights into the underlying mechanisms.
@@ 2.1.1 Covered setup: Circular waves initiate synchronously at the pillars and maintain their circular shape
Figure 2: Number of petals depends on the pillar diameter and can vary over time. (A) A PDMS mold with alternating pillar diameters of 1 mm and 1.5 mm is shown. (B) A sketch illustrates the covered versus uncovered experimental setup. (C) In the uncovered setup with evaporative cooling, the evaporation rate is approximately 2.5$\times$10-3 mg/(min mm2), which is about seven times higher than in the covered setup. (D-E) Top views of chemical waves in an uncovered setup captured before and after the onset of instability in an experiment where pillar diameters alternate between rows, with 1.5 mm pillars in the top row and 1 mm pillars in the second row. Pillars with a diameter of 1.5 mm generally develop four petals, whereas those with a diameter of 1 mm typically produce three petals. (F-H) Space-time plots along the highlighted circles in panel (D) illustrate the progression of the instability. (I) An example of a pillar that initially developed three petals, but over time, two of the petals merged.
In an uncovered setup, where evaporation is more significant (Figure 2C), the obstacles continue to act as centers that emit synchronous concentric waves (Figure 1E, I and Video 3). However, after approximately two cycles of these circular waves, an instability in the wavefront becomes apparent: the circular wavefronts fragment, leading to the formation of striking flower-like patterns (see Figure 1F, J). Figure 1L and Video 4 present a separate experiment where the entire Petri dish is imaged, demonstrating similar wavefront instability. Consistent with the covered setup, Fourier analysis shows that the waves originating from the obstacles typically have a periodicity of about 45 seconds (see Figure S1E-H).
@@
Figure 3: Wavefront instability at the boundary in an uncovered setup. (A-B) Pillars with a diameter of 1.85 mm typically form four petals. The color-coded phase map, generated using Hilbert transform, indicates that the pillars simultaneously act as wave centers. (C) A space-time plot along the cyan-highlighted circle in panel (A) shows the development of the instability, with a noticeable evolution in the number of petals over time. The wave period is 50 sec. (D) A space-time plot along the orange-highlighted circle demonstrates the wavefront instability that also manifests at the PDMS wall on the periphery. This experiment is conducted in a small Petri dish.
Figure 4: Effect of the pillar diameter on the number of petals. (A-D) As the pillar diameter decreases, so does the number of petals formed. For instance, pillars with a diameter of 3.8 mm typically develop around 7 petals, while those with a diameter of 2.8 mm produce about 6 petals. This number drops to 4 petals for pillars with a diameter of 1.8 mm, and further reduces to 3 petals for those with a diameter of 1 mm. (E) Pillars with a diameter of 0.5 mm exhibit no wavefront instability, which verifies that there is a minimum pillar perimeter required for such instability to develop. (F) The number of petals increases linearly with the diameter of the pillar. Below a critical diameter, wavefronts do not break and remain circular.
@@
Figure 6: Hydrophilic versus hydrophobic obstacles: An overview of a partitioned, uncovered Petri dish taken at two distinct times: before (A) and after (B) the emergence of instability. The upper portion of the dish has been plasma-treated to render PDMS hydrophilic, whereas the lower section has been left untreated, maintaining its hydrophobic properties (see Video 9). A zoomed-in side view of the fluid near the pillars is also presented, emphasizing the rising fluid around the hydrophilic pillar. (C-E) Side-view imaging of the fluid climbing the hydrophilic pillars confirmed that the contact line remains pinned. As the liquid evaporates, the contact angle gradually decreases (see Video 12). The cyan curves represent the air-liquid interface at $t$$=$0.
@@ 2.1.3 Experimental evidence of a fingering instability
Moreover, side-view imaging of the fluid climbing around the pillars confirms that the contact line remains pinned (Figure 6C-E and Video 12), indicating the presence of upward capillary flows that counteract fluid loss caused by evaporation. Within this ascending liquid layer, the thinner downstream precursor film is expected to evaporate faster than the thicker upstream film, creating slight temperature and/or concentration gradients. These gradients generate surface tension differences, driving Marangoni flows that ultimately disrupt the wavefront. Our hypothesis that waves originate from the thin ascending fluid layer is supported by our observations in Figure 5, using a PMMA mold with square wall boundaries. The side view of the waves confirms that they initiate at the thin film in the contact region, which wets the square wall boundaries, and then travel downward (Video 13). Importantly, in an uncovered setup, the downward-traveling wave along the wall boundary (as well as the waves initiated around the cylindrical obstacles) initially maintain a planar (circular) form but eventually break into fingers at later times (Figure 5B-C and D). This fingering instability is similarly visible in Figure 8E and Video 14 with PMMA-based triangular obstacles within square wall boundaries. Lastly, we note that the initiation of waves at the ascending thin film is consistent with a separate experiment in which the Petri dish was slightly tilted [37]. In this setup, we observed that synchronization waves initially appeared in regions with lower fluid thickness and subsequently propagated toward areas with higher fluid thickness (see Figure S5A-B and Video 15).
@@
Literature reveals that Marangoni-driven thin films on planar or curved substrates exhibit a complex array of dynamics [57, 54, 69], dependent on the flat upstream film thickness trailing the advancing front. When the upstream film thickness is set by the competition effect of meniscus curvature and the surface tension gradient, the film is sufficiently thin for the advancing front to form a stationary compressive or Lax shock, which are well known to be unstable to contact line perturbations [69, 54, 57, 55]. It is important to note that surface tension gradients, which drive Marangoni flows, can result from temperature gradients, concentration gradients, or a combination of both. Concentration gradients often develop due to the evaporation of more volatile chemical components within the ascending thin film, similar to the mechanism observed in the tears of wine phenomenon [70]. The interplay between Marangoni flows and gravity can disrupt the contact lines, evolving into a finger pattern. Specifically, the linear stability analysis in Ref. [57] correlates linearly the number of fingers with the cylindrical pillar’s perimeter, aligning with our findings (see Figure 4). This analysis also suggests a minimum pillar diameter threshold below which fingering instability does not occur. To validate this, we reduced the pillar diameter to 0.5 mm in our uncovered setup experiments, observing that the wavefronts remained nearly circular (see Figure 4E and Figure S3).
@@
Figure 9: Experiments with both inverted tapered and tapered pillars. (A) Conical pillars with base diameters ranging from 0.5 mm to 2 mm, with the top row displaying standard cylindrical pillars of 3 mm diameter. (B) Side view of centered waves observed before and after the onset of instability. The space-time within the green box represents the light intensity along the green line over time, revealing that the waves originate from the climbing thin film and propagate downward. (C-D) A close-up side view of two neighboring cones with a base diameter of 1 mm. The cyan curve illustrates the air-liquid interface at $t$$=$0. (E) Top view of the centered waves prior to the instability. (F) Tapered pillars with a 3 mm base diameter that gradually decrease to smaller diameters toward the top. (G-J) The side and top views of the centered waves are shown, along with the contact angle of the CHD-BZ solution around the hydrophilic conical structures at $t$$=$0. Over time, as evaporation progresses, the contact angle decreases ($t$$=$300 sec). The cyan curves in panel (H) represent the liquid-air interface at $t$$=$0; see Videos 17-20.
@@
Furthermore, we observed that during multiple cycles of covering and uncovering the setup, the wavefronts, which exhibit flower-like patterns when uncovered, revert to their circular form once the setup is covered, and vice versa (Videos 21-22). To further investigate the role of boundaries, we conducted a control experiment using an uncovered, plasma-treated glass Petri dish without any obstacles (Figure S4 and Video 23). Interestingly, we observed wavefront instability at the glass boundary, closely resembling the phenomena observed at the PDMS wall in Figure 3 and the PMMA wall in Figures 5 and 8E. A similar control experiment was performed in an uncovered Petri dish with a PDMS substrate (no pillars) and a PDMS wall boundary, where wavefront instability was only observed at the PDMS wall (see Video 24). Overall, these experiments confirm that the thin layer of climbing fluid at the boundary is critical for initiating waves, and that evaporation significantly contributes to the wavefront instability.
@@
Figure 11: Symmetry breaking in an evaporating CHD-BZ droplet and coffee stain effect. (A-D) A droplet of volume $2.5~{}\mu l$ is deposited on a hydrophobic glass substrate. After about 500 sec, the azimuthal symmetry breaks and spots form at the rim evolving over time ($T$$=$22C, RH$=$$40\%$, $\theta_{0}$$=$12). The images are captured in the reflection mode of a confocal microscope; see Video 25. (E-H) In a separate experiment imaging the entire droplet, micron-sized beads were introduced to track the flow. During the first 5 minutes of evaporation, the droplet boundary remains pinned to the substrate, leading to the accumulation of tracer particles at the rim due to capillary flows, clearly exhibiting the coffee-stain effect, as depicted in panels (F-H). As evaporation continues and the reaction progresses, a secondary instability develops at the rim of the shrinking droplet, as shown in panel (H) (see Video 26).

### v4

@@ Abstract
Far from equilibrium, chemical and biological systems can form complex patterns and waves through reaction-diffusion coupling. Fluid motion often interferes with these self-organized concentration patterns. In this study, we investigate the influence of Marangoni-driven flows inside a thin layer of fluid ascending the outer surfaces of hydrophilic obstacles on the spatio-temporal dynamics of chemical waves in the modified Belousov-Zhabotinsky reaction. Our observations reveal that circular waves originate nearly simultaneously at the obstacles and propagate outward. In a covered setup, where evaporation is minimal, the wavefronts maintain their circular shape. However, in an uncovered setup with significant evaporation and resulting Marangoni flows, the interplay between surface tension-driven Marangoni flows and gravity destabilizes the wavefronts, creating distinctive flower-like patterns around the obstacles. Our analysis shows that here solutal Marangoni forces are more relevant than thermal ones. Our experiments further show that the number of petals formed increases linearly with the obstacle’s diameter, though a minimum diameter is required for these instabilities to appear. These findings demonstrate the potential to ’engineer’ specific wave patterns, offering a method to control and direct reaction dynamics. This capability is especially important for developing microfluidic devices requiring precise control over chemical wave propagation.
@@
Sangram Gore, Binaya Paudyal, Duarte Rocha, Mohamed Ali, Nader Masmoudi, Albert Bae, Christian Diddens, Detlef Lohse, Oliver Steinbock, Azam Gholami*
@@
D. Rocha, Dr. C. Diddens, Prof. D. Lohse
Physics of Fluids Department, Max-Planck Center Twente for Complex Fluid Dynamics and J. M. Burgers Centre for Fluid Dynamics, University of Twente, Enschede, The Netherlands
Prof. D. Lohse
Max-Planck Institute for Dynamics and Self-Organization, Am Faßberg 17, 37077 Göttingen, Germany
@@
Lewis &\& Clark College, Portland, Oregon, USA
@@ 1 Introduction
The interaction between chemical reactions and convective fluid motion, particularly through mechanisms such as Marangoni-driven flows, profoundly influences the formation and breakdown of complex spatio-temporal patterns [26, 27, 28, 29, 30]. In pioneering studies on the BZ reaction [31, 32, 33, 34, 35], it was observed that in confined environments, chemical instabilities manifest themselves as propagating waves. In contrast, in systems with a free surface, the presence of Marangoni flows lead instead to the formation of cellular structures [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 12]. These flows are initiated by gradients in surface tension [47, 45], which might arise from either temperature variations or differences in solute concentrations across the fluid’s surface, causing the fluid to flow from regions of lower to higher surface tension. For instance, in the ferroin-catalyzed BZ reaction, the Fe2+ complex has a higher surface activity than the Fe3+ complex which is attributed to the differing hydrophilic properties of the catalyst’s oxidized (more hydrophilic) and reduced forms. This difference leads to a decrease in surface tension of approximately 2 mN/m when the system transitions from the reduced to the oxidized state [47]. In an uncovered setup, where evaporation is significant, the interplay of chemical waves, buoyancy-driven flows due to density gradients, and thermal and solutal Marangoni flows due to surface tension gradients collectively influences the resulting chemo-hydrodynamic patterns [32, 48, 49, 42, 41, 50, 35, 30].
@@
Evaporation-induced Marangoni flows can be intensified by introducing hydrophilic vertical or inclined obstacles within the reactive medium. The thin film rising along the outer surfaces of these obstacles evaporates more quickly than the bulk fluid, creating thermal and/or solutal surface tension gradients by evaporative cooling and/or different volatilities of the multicomponent liquid. The delicate balance between upward-driving surface tension forces and gravity-induced downward motion leads to undulations at the fluid interface, potentially developing into finger-like patterns if the fluid becomes sufficiently unstable [51, 52, 53, 54, 55, 56, 57]. The obstacle’s geometry also affects the dynamics at the droplet-gas of thin films [57, 58]. A cylindrical obstacle introduces unique conditions, such as azimuthal curvature and periodic boundary conditions, absent in flat obstacles [57]. This periodicity ensures that an integer number of fingers can form along a cylinder, a condition that is not required for a flat obstacle, and the smallest number of fingers that can form along the contact line of a cylinder is one [57].
@@
Building on these groundworks, our study introduces a novel approach to explore the intricate interaction between evaporation-driven Marangoni flows and chemical waves in the modified BZ reaction around hydrophilic obstacles. In a covered setup – where evaporation and therefore the evaporation-induced surface tension differences are minimal – these obstacles act as focal points, emitting synchronous circular waves that retain their shape over multiple cycles. By contrast, in an uncovered setup with pronounced evaporation, thermal and solutal Marangoni flows trigger a hydrodynamic instability, yielding characteristic flower-like wave patterns. The number of ‘petals’ scales with obstacle diameter, and a critical diameter threshold is required for the instability to emerge. We also show that in our system, solutal Marangoni forces are more relevant than thermal ones. To elucidate the driving mechanism, we perform numerical simulations of a simplified binary fluid (e.g., water–ethanol) subject to solutal Marangoni instabilities, due to selective evaporation of one component (e.g. ethanol as compared to water). These simulations qualitatively reproduce the observed symmetry-breaking patterns and lay the groundwork for future quantitative modeling.
@@ 2.1.1 Covered setup: Circular waves initiate synchronously at the pillars and maintain their circular shape
[…]
In an uncovered setup, where evaporation is more significant (Figure 2C), the obstacles continue to act as centers that emit synchronous concentric waves (Figure 1E, I and Video 3). However, after approximately two cycles of these circular waves, an instability in the wavefront becomes apparent: the circular wavefronts fragment, leading to the formation of striking flower-like patterns (see Figure 1F, J). Figure 1L and Video 4 present a separate experiment where the entire Petri dish is imaged, demonstrating similar wavefront instabilities. Consistent with the covered setup, Fourier analysis shows that the waves originating from the obstacles typically have a periodicity of about 45 seconds (see Figure S1E-H).
@@
Figure 3: Wavefront instability at the boundary in an uncovered setup. (A-B) Pillars with a diameter of 1.85 mm typically form four petals. The color-coded phase map, generated using Hilbert transform, indicates that the pillars simultaneously act as wave centers. (C) A space-time plot along the cyan-highlighted circle in panel (A) shows the development of the instability, with a noticeable evolution in the number of petals over time. The wave period is 50 sec. On the right, these snapshots at three different times ($126\text{\,}\mathrm{s}$, $180\text{\,}\mathrm{s}$ and $233\text{\,}\mathrm{s}$) are shown. (D) A space-time plot along the orange-highlighted circle demonstrates the wavefront instability that also manifests at the PDMS wall on the periphery. This experiment is conducted in a small Petri dish.
Figure 4: Effect of the pillar diameter on the number of petals. (A-D) As the pillar diameter decreases, so does the number of petals formed. For instance, pillars with a diameter of 3.8 mm typically develop around 7 petals, while those with a diameter of 2.8 mm produce about 6 petals. This number drops to 4 petals for pillars with a diameter of 1.8 mm, and further reduces to 3 petals for those with a diameter of 1 mm. (E) Pillars with a diameter of 0.5 mm exhibit no wavefront instability, which verifies that there is a minimum pillar perimeter required for such instability to develop. (F) The number $N$ of petals increases linearly with the diameter $d$ of the pillar. Below a critical diameter, wavefronts do not break and remain circular. The inset shows the average lateral extension $\pi d/N$ of the petals.
@@
Figure 6: Hydrophilic versus hydrophobic obstacles: An overview of a partitioned, uncovered Petri dish taken at two distinct times: before (A) and after (B) the emergence of instability. The upper portion of the dish has been plasma-treated to render PDMS hydrophilic, whereas the lower section has been left untreated, maintaining its hydrophobic properties (see Video 9). A zoomed-in side view of the fluid near the pillars is also presented, emphasizing the rising fluid around the hydrophilic pillar. (C-E) Side-view imaging of the fluid climbing the hydrophilic pillars confirmed that the contact line remains pinned. As the liquid evaporates, the contact angle gradually decreases (see Video 12). The cyan curves represent the air-liquid interface at tt==0; See the SI for meniscus analysis and curvature calculations.
@@ 2.1.3 Experimental evidence of a fingering instability
Moreover, side-view imaging of the fluid climbing around the pillars confirms that the contact line remains pinned (Figure 6C-E and Video 12), indicating the presence of upward capillary flows that counteract fluid loss caused by evaporation. Within this ascending liquid layer, the thinner downstream precursor film is expected to evaporate faster than the thicker upstream film, creating slight temperature and/or concentration gradients, due to evaporative cooling and/or selective evaporation of one or more components of the multicomponent liquid. These gradients generate surface tension differences, driving thermal and/or solutal Marangoni flows that ultimately disrupt the wavefront. Our hypothesis that waves originate from the thin ascending fluid layer is supported by our observations in Figure 5, using a PMMA mold with square wall boundaries. The side view of the waves confirms that they initiate at the thin film in the contact region, which wets the square wall boundaries, and then travel downward (Video 13). Importantly, in an uncovered setup, the downward-traveling wave along the wall boundary (as well as the waves initiated around the cylindrical obstacles) initially maintain a planar (circular) form but eventually break into fingers at later times (Figure 5B-C and D). This fingering instability is similarly visible in Figure 8E and Video 14 with PMMA-based triangular obstacles within square wall boundaries. Lastly, we note that the initiation of waves at the ascending thin film is consistent with a separate experiment in which the Petri dish was slightly tilted [38]. In this setup, we observed that synchronization waves initially appeared in regions with lower fluid thickness and subsequently propagated toward areas with higher fluid thickness (see Figure S5A-B and Video 15).
@@
Literature reveals that Marangoni-driven thin films on planar or curved substrates exhibit a complex array of dynamics [57, 54, 70, 71], dependent on the flat upstream film thickness trailing the advancing front. When the upstream film thickness is set by the competition effect of meniscus curvature and the surface tension gradient, the film is sufficiently thin for the advancing front to form a stationary compressive or Lax shock, which are well known to be unstable to contact line perturbations [70, 54, 57, 55]. It is important to note that surface tension gradients, which drive Marangoni flows, can result from temperature gradients, concentration gradients, or a combination of both. Concentration gradients often develop due to the evaporation of more volatile chemical components within the ascending thin film, similar to the mechanism observed in the tears of wine phenomenon [72, 73]. The interplay between Marangoni flows and gravity can disrupt the contact lines, evolving into a finger pattern. Specifically, the linear stability analysis in Ref. [57] correlates linearly the number of fingers with the cylindrical pillar’s perimeter, aligning with our findings (see Figure 4). This analysis also suggests a minimum pillar diameter threshold below which fingering instability does not occur. To validate this, we reduced the pillar diameter to 0.5 mm in our uncovered setup experiments, observing that the wavefronts remained nearly circular (see Figure 4E and Figure S3). This resembles the situation of Rayleigh-Bénard convection where, for given driving strength, the linear instability cannot develop for containers with a too small aspect ratio (width/height) [74].
@@
Figure 9: Experiments with both inverted tapered and tapered pillars. (A) Conical pillars with base diameters ranging from 0.5 mm to 2 mm, with the top row displaying standard cylindrical pillars of 3 mm diameter. (B) Side view of centered waves observed before and after the onset of instability. The space-time within the green box represents the light intensity along the green line over time, revealing that the waves originate from the climbing thin film and propagate downward. (C-D) A close-up side view of two neighboring cones with a base diameter of 1 mm. The cyan curve illustrates the air-liquid interface at tt==0. (E) Top view of the centered waves prior to the instability. (F) Tapered pillars with a 3 mm base diameter that gradually decrease to smaller diameters toward the top. (G-J) The side and top views of the centered waves are shown, along with the contact angle of the CHD-BZ solution around the hydrophilic conical structures at tt==0. Over time, as evaporation progresses, the contact angle decreases (tt==300 sec). The cyan curves in panel (H) represent the liquid-air interface at tt==0; see Videos 17-20.
@@
Furthermore, we observed that during multiple cycles of covering and uncovering the setup, the wavefronts, which exhibit flower-like patterns when uncovered, revert to their circular form once the setup is covered, and vice versa (Videos 21-22). To further investigate the role of boundaries, we conducted a control experiment using an uncovered, plasma-treated glass Petri dish without any obstacles (Figure S4 and Video 23). Interestingly, we observed wavefront instability at the glass boundary, closely resembling the phenomena observed at the PDMS wall in Figure 3 and the PMMA wall in Figures 5 and 8E. A similar control experiment was performed in an uncovered Petri dish with a PDMS substrate (no pillars) and a PDMS wall boundary, where the wavefront instability was only observed at the PDMS wall (see Video 24). Overall, these experiments confirm that the thin layer of climbing fluid at the boundary is critical for initiating the waves, and that evaporation significantly contributes to the wavefront instability.
@@
[…]